OpenStudy (anonymous):

what is the LCM of 96 and 212?

OpenStudy (anonymous):

96 : 2 = 48 :2 = 24 : 2 = 12 : 2 =6: 2 =3 96 = 2^5 *3 212 : 2 = 106 : 2 = 53 212 = 2^2 * 53 lcm = 2^5 * 53 = 1696